OpenStudy (anonymous):

PLEASE HELP! Grade 9 Question: Find the root of each equation using pencil and paper. Check each root. a.) 1/2 (x+3)=5 b.) b-5/7 =3 c.) 6= 2/3 m-1 d.) -5= 3d+4/3

OpenStudy (anonymous):

please don't tell me the answers, I just want to know how to do it, step by step, because I have a math exam tomorrow :)

OpenStudy (anonymous):

a.) 1/2 (x+3)=5 multiply both sides by 2 x+3 = 10 now subtract 3 from both sides x= 7

OpenStudy (anonymous):

you understand what i did??

OpenStudy (anonymous):

hi sorry

OpenStudy (anonymous):

in each equation you have one variable now its you job to separate that variable one one side and put all the constants on other side using maths

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what happens to the fraction?

OpenStudy (anonymous):

look at you question carefully a.) 1/2 (x+3)=5

OpenStudy (anonymous):

you have to find the root for this equation. root means to find such value of x that satisfies this eqquation

OpenStudy (anonymous):

oh...i see now

OpenStudy (anonymous):

how would you do d?

OpenStudy (anonymous):

a.) 1/2 (x+3)=5 now if you multiply both sides by 2 what you get ????

OpenStudy (anonymous):

1(x+3)=10

OpenStudy (anonymous):

10-3

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

exactly you got x= 7 which is the required root for this equation

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

to verify this that either your root is correct or incorrect just plug in this value in x in you question and see whether you get LHS = RHS or not if LHS = RHS then you calculated the correct root else wrong value

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

can you do the rest of questions or not???

OpenStudy (anonymous):

yea, i think so, thanks

OpenStudy (anonymous):

welcomeeeeeeee

OpenStudy (anonymous):

:)